# Restore-RWKV: Efficient and Effective Medical Image Restoration with RWKV

### v2

@@ Abstract
Transformers have revolutionized medical image restoration, but the quadratic complexity still poses limitations for their application to high-resolution medical images. The recent advent of RWKV in the NLP field has attracted much attention as it can process long sequences efficiently. To leverage its advanced design, we propose Restore-RWKV, the first RWKV-based model for medical image restoration. Since the original RWKV model is designed for 1D sequences, we make two necessary modifications for modeling spatial relations in 2D images. First, we present a recurrent WKV (Re-WKV) attention mechanism that captures global dependencies with linear computational complexity. Re-WKV incorporates bidirectional attention as basic for a global receptive field and recurrent attention to effectively model 2D dependencies from various scan directions. Second, we develop an omnidirectional token shift (Omni-Shift) layer that enhances local dependencies by shifting tokens from all directions and across a wide context range. These adaptations make the proposed Restore-RWKV an efficient and effective model for medical image restoration. Extensive experiments demonstrate that Restore-RWKV achieves superior performance across various medical image restoration tasks, including MRI image super-resolution, CT image denoising, PET image synthesis, and all-in-one medical image restoration. Code is available at: https://github.com/Yaziwel/Restore-RWKV.
@@ 1 Introduction
Medical image restoration (MedIR) aims at recovering the high-quality (HQ) medical image from its degraded low-quality (LQ) counterpart. It encompasses a variety of tasks such as MRI image super-resolution [1, 2, 3, 4, 5, 6, 7], CT image denoising [8, 9, 10], PET image synthesis [11, 12, 13, 14, 15], and all-in-one medical image restoration [16, 17, 18, 19]. This field is particularly challenging due to the ill-posed nature of the problem, where crucial information about the image content is often lost in LQ images. Recent years have witnessed significant progress in MedIR, largely driven by advanced deep learning models such as convolutional neural networks (CNNs) [20, 21], Transformers [22, 23], and Mambas [24, 25, 26].
@@
Figure 1: The effective receptive field (ERF) [27] visualization for different models. A more extensively distributed dark area indicates a larger ERF. Our proposed Restore-RWKV achieves the most significant global ERF.
@@
In spite of significant advancements, current CNN-based, Transformer-based, and Mamba-based models still exhibit inherent limitations for accurate medical image restoration. CNN-based models [2] often face constraints in their effective receptive field [27] (ERF, as illustrated in Figure 1 (a)) due to the limited kernel size of the convolution operator. This limitation hampers their ability to capture information across broader ranges to compensate for lost details in degraded pixels. In contrast, Transformer-based models excel in modeling long-range dependencies and achieving a global ERF through self-attention [23]. However, the computational complexity of self-attention grows quadratically with spatial resolution, rendering standard Transformers impractical for high-resolution medical images. To mitigate this, many other Transformer-based models aim to reduce computational demands by restricting attention to smaller windows [28], inevitably diminishing the ERF (as depicted in Figure 1 (b)) and conflicting with the goal of capturing long-range dependencies. Mamba-based models [29, 7] serve as an efficient alternative to Transformers, leveraging state space models (SSMs) [24] to causally capture long-range dependencies with linear computational complexity. Recent studies suggest that Mamba-based models can achieve performance comparable to or even better than Transformer-based models. However, their inherent unidirectional sequence modeling in SSMs still poses challenges in achieving an optimal ERF for 2D images (as shown in Figure 1 (c)). In summary, existing models in MedIR struggle to strike a good balance between computational efficiency and achieving an effective global receptive field.
@@
In this paper, we propose Restore-RWKV, an efficient and effective model adapted from RWKV for medical image restoration. The original RWKV [30] is designed for handling 1D sequences and has limitations in capturing spatial dependencies in 2D medical images. To address these limitations, we make essential modifications to two key components of RWKV—the attention layer and the token shift layer—to adapt it for processing 2D medical images. (1) We introduce a recurrent WKV (Re-WKV) attention mechanism that effectively captures global dependencies with linear computational complexity. Unlike the causal WKV attention in the original RWKV, which is unidirectional and has a limited receptive field, Re-WKV utilizes a bidirectional attention mechanism to achieve a global receptive field. Additionally, it employs a recurrent attention mechanism that effectively models 2D dependencies across various scan directions. (2) We develop an omnidirectional token shift (Omni-Shift) layer to accurately capture local dependencies. Previous token shift layers only shift adjacent tokens from limited directions using simple interpolation. In contrast, Omni-Shift enhances interaction among neighbors by shifting tokens from all directions and over a large context range through efficient depth-wise convolution. Furthermore, during training, Omni-Shift employs a structural re-parameterization strategy to learn to accurately shift tokens from various context ranges, while maintaining the original structure for efficiency during testing. Extensive experimental results indicate that our proposed Restore-RWKV can achieve superior performance with good efficiency, serving as a general restoration backbone for various tasks, including MRI image super-resolution, CT image denoising, PET image synthesis, and all-in-one medical image restoration.
@@ 2.1 Medical Image Restoration
Medical image restoration (MedIR) aims to recover high-quality (HQ) medical images from their degraded low-quality (LQ) counterparts. Typical MedIR tasks include MRI image super-resolution [1, 2, 3, 4, 5, 6, 7], CT image denoising [8, 9, 10], PET image synthesis [11, 12, 13, 14, 15], and all-in-one medical image restoration for multi-task MedIR [16, 17, 18, 19]. Recently, deep learning-based models have become the primary approach for addressing MedIR tasks. These methods can be broadly categorized into CNN-based models, Transformer-based models, and Mamba-based models.
CNN-based Models. Over the past decades, CNN-based models [4, 5, 8, 11, 13] have made extraordinary contributions to the field of MedIR due to their strong representation capabilities. A key advantage of CNN-based models for MedIR is their ability to capture local context and build local dependencies through convolution. This allows the neighborhood of a degraded pixel to be used as a reference to aid in its recovery. However, due to the limited kernel size of convolutions, CNN-based methods often suffer from restricted receptive fields and cannot model long-range dependencies. While some approaches have tried to expand the receptive field by designing deeper networks [2, 3] or utilizing the multi-scale U-Net [8, 13] architecture, these improvements are still limited and struggle to capture long-range dependencies effectively.
@@
Mamba-based Models. Mamba [24, 26, 25], a sequence model backbone grounded in state-space models (SSMs), has emerged as a prominent alternative to Transformers. Unlike Transformers, which exhibit quadratic computational complexity, Mamba-based models can model long-range dependencies with linear computational complexity. This capability allows Mamba-based models to directly capture global dependencies in high-resolution medical images through SSMs without the need for efficient designs like window partitioning. Recent studies have shown that Mamba-based models [29, 7] outperform Transformer-based models in MedIR tasks. However, due to the unidirectional nature of SSM, Mamba’s global dependency modeling capability still has limitations in achieving accurate results for MedIR applications.
@@ 3 Method
We aim to build an efficient and effective model for medical image restoration based on the Receptance Weighted Key Value (RWKV) [30], termed Restore-RWKV. In Sec. 3.1, we introduce the network architecture of Restore-RWKV, and in Sec. 3.2, we present a novel R-RWKV block as its fundamental component for feature extraction. Since the original RWKV [30] is designed for processing 1D sequences, we incorporate two innovations in the R-RWKV block to adapt it for capturing global and local dependencies in 2D images. First, in Sec. 3.3, we introduce a recurrent WKV (Re-WKV) attention mechanism that effectively builds global dependencies in 2D images with linear complexity. Second, in Sec. 3.4, we introduce an omnidirectional token shift that accurately captures local context to enhance the model performance. We will elaborate on the details of Restore-RWKV in the following section.
@@ 3.1 Restore-RWKV Architecture
As shown in Figure 2 (a), the proposed Restore-RWKV is a 4-level U-shaped encoder-decoder architecture with 3 times downsampling and upsampling, which enjoys the advantage of capturing image features at different hierarchies and computational efficiency [39]. Given an input low-quality (LQ) image $I_{LQ}\in\mathbb{R}^{H\times W\times 1}$, Restore-RWKV first employs a $3\times 3$ convolutional layer as an input projection to project the input image to a shallow feature $F_{S}\in\mathbb{R}^{H\times W\times C}$, where $H$, $W$, and $C$ denote the height, width, and channel, respectively. Then the shallow feature $F_{S}$ undergoes a 4-level hierarchical encoder-decoder and is transformed into a deep feature $F_{D}\in\mathbb{R}^{H\times W\times 2C}$. Each level of the encoder-decoder comprises $N_{i}$, where $i\in\{1,2,3,4\}$, R-RWKV blocks for feature extraction, with an additional $N_{\text{refinement}}$ R-RWKV blocks used for feature refinement in the final decoder level. For feature downsampling, Restore-RWKV utilizes a $1\times 1$ convolutional layer and a pixel-unshuffle operation [40], reducing the spatial size by half and doubling the channel number. For feature upsampling, Restore-RWKV employs a pixel-shuffle operation [40] and a $1\times 1$ convolutional layer, doubling the spatial size and halving the channel number. To aid in the restoration process, encoder features are concatenated with decoder features via skip connections. The deep feature $F_{D}$ is then projected out to a residual image $I_{R}\in\mathbb{R}^{H\times W\times 1}$ by a $3\times 3$ convolutional layer. Finally, the restored image $\hat{I}_{HQ}$ can be obtained by adding the input $I_{LQ}$ and the residual image $I_{R}$: $\hat{I}_{HQ}=I_{LQ}+I_{R}$.
@@ 3.3 Recurrent WKV Attention
Here, $\operatorname{Bi-WKV}^{(j)}(\cdot)$ denotes the $j$-th Bi-WKV attention. $\operatorname{\Delta}_{dir}(\cdot)$ represents the changing direction operation. In two adjacent Bi-WKV attention operations, we alternately use two different scan directions: horizontal scan (H-Scan) and vertical scan (V-Scan). $wkv^{(j)}$ and $wkv^{(j-1)}$ are attention results of the $j$-th and $(j-1)$-th Bi-WKV, respectively. Note that $wkv^{(0)}$ is the input $V_{s}$. Eq. 9 indicates the current $j$-th Bi-WKV attention takes the previous $(j-1)$-th Bi-WKV attention result as its input value, complementing the $(j-1)$-th attention result from a different scan direction. The final attention result $wkv$ is obtained after applying Bi-WKV attention recurrently $M$ times:
@@
In comparison with Bi-WKV, Re-WKV enhances global token interactions in 2D images by recurrently performing attention along different scan directions. Furthermore, since $M\ll T$ in our implementation, the computational complexity of Re-WKV maintains the linear computational complexity as Bi-WKV.
@@ 3.4 Omnidirectional Token Shift
where $\alpha_{i}$ denotes a learnable parameter for scaling the specific branch. $\operatorname{DConv}_{k\times k}(\cdot)$ denotes the efficient depth-wise convolution with kernel size of $k$. The final result of Omni-Shift is a fusion of four branches, with each branch specializing in a specific context range, resulting in an accurate token shift result.
@@ 4.1 Dataset
We conduct extensive experiments on various MedIR tasks to demonstrate the performance of our proposed Restore-RWKV, including MRI image super-resolution, CT image denoising, PET image synthesis, and an all-in-one medical image restoration task that combines the aforementioned three tasks for multi-task learning. The corresponding datasets for each task are introduced as follows.
@@
CT Image Denoising. We utilize the publicly available AAPM dataset  [^1] https://www.aapm.org/GrandChallenge/LowDoseCT/, which comprises paired standard-dose HQ CT images and quarter-dose LQ CT images. These images are collected from 10 patients and we divide them into 8 patients for training, 1 for validation, and 1 for testing. Each 3D CT image has a dimension of 512$\times$512×$\times n$. Consequently, we extract 2D slices sized 512$\times$512 and obtain 2039 images for training, 128 images for validation, and 211 images for testing.
PET Image Synthesis. We collect HQ PET images from 159 patients using the PolarStar m660 PET/CT system in list mode, with an injection dose of 293MBq 18F-FDG. These 159 HQ PET images were divided into 120 for training, 10 for validation, and 29 for testing. LQ PET images are generated through random list mode decimation with a dose reduction factor of 12. Both HQ and LQ PET images are reconstructed using the standard OSEM [42] method. Each PET image has 3D shapes of 192$\times$192$\times$400 and is divided into 192 2D slices sized 192$\times$400. Then, slices containing only air are excluded. Consequently, we obtain 8350 2D PET images for training, 684 images for validation, and 2044 images for testing.
All-in-One Medical Image Restoration. The goal of all-in-one medical image restoration [18] is to train a universal model capable of handling multiple MedIR tasks. Therefore, the dataset we use is a combination of the datasets from the aforementioned MRI image super-resolution, CT image denoising, and PET image synthesis tasks.

### v3

@@ Abstract
Transformers have revolutionized medical image restoration, but the quadratic complexity still poses limitations for their application to high-resolution medical images. The recent advent of the Receptance Weighted Key Value (RWKV) model in the natural language processing field has attracted much attention due to its ability to process long sequences efficiently. To leverage its advanced design, we propose Restore-RWKV, the first RWKV-based model for medical image restoration. Since the original RWKV model is designed for 1D sequences, we make two necessary modifications for modeling spatial relations in 2D medical images. First, we present a recurrent WKV (Re-WKV) attention mechanism that captures global dependencies with linear computational complexity. Re-WKV incorporates bidirectional attention as basic for a global receptive field and recurrent attention to effectively model 2D dependencies from various scan directions. Second, we develop an omnidirectional token shift (Omni-Shift) layer that enhances local dependencies by shifting tokens from all directions and across a wide context range. These adaptations make the proposed Restore-RWKV an efficient and effective model for medical image restoration. Even a lightweight variant of Restore-RWKV, with only 1.16 million parameters, achieves comparable or even superior results compared to existing state-of-the-art (SOTA) methods. Extensive experiments demonstrate that the resulting Restore-RWKV achieves SOTA performance across a range of medical image restoration tasks, including PET image synthesis, CT image denoising, MRI image super-resolution, and all-in-one medical image restoration. Code is available at: https://github.com/Yaziwel/Restore-RWKV.
@@ 1 Introduction
Medical image restoration (MedIR) aims at recovering the high-quality (HQ) medical image from its degraded low-quality (LQ) counterpart. It encompasses a variety of tasks such as PET image synthesis [1, 2, 3, 4, 5, 6, 7, 8], CT image denoising [9, 10, 11, 12, 13, 14, 15], MRI image super-resolution [16, 17, 18, 19, 20], and all-in-one medical image restoration [21, 8, 22, 23]. This field is particularly challenging due to the ill-posed nature of the problem, where crucial information about the image content is often lost in LQ images. Recent years have witnessed significant progress in MedIR, largely driven by advanced deep learning models such as convolutional neural networks (CNNs) [24, 25], Transformers [26, 27], and Mambas [28, 29, 15, 30].
@@
Figure 1: The effective receptive field (ERF) [31] visualization for different efficient models. A more extensively distributed dark area indicates a larger ERF. Our proposed Restore-RWKV achieves the most significant global ERF.
@@
Despite significant advancements, current CNN-based, Transformer-based, and Mamba-based models face challenges in striking a proper balance between a large receptive field and computational efficiency. To some extent, improvements in medical image restoration performance are largely dependent on the model’s receptive field. On one hand, a larger receptive field enables the model to capture information from a wider region, allowing it to reference more pixels to improve the recovery of the degraded areas. On the other hand, a larger receptive field allows the model to capture high-level patterns and structures, which are essential for effective medical image restoration [32]. However, increasing the receptive field often leads to higher computational complexity in model design. Among existing studies, CNN-based models [33] often face constraints in their effective receptive field [31] (ERF, as illustrated in Fig. 1 (a)) due to the limited kernel size of the convolution operator. This limitation hampers their ability to capture information across broader ranges to compensate for lost information in degraded pixels. In contrast, Transformer-based models excel in modeling long-range dependencies and achieving a global ERF through self-attention [27]. However, the computational complexity of self-attention grows quadratically with spatial resolution, rendering standard Transformers impractical for high-resolution medical images. To mitigate this, many other Transformer-based models aim to reduce computational demands by restricting attention to smaller windows [18], inevitably diminishing the ERF (as depicted in Fig. 1 (b)) and conflicting with the goal of capturing long-range dependencies. Mamba-based models [32, 34, 15] serve as an efficient alternative to Transformers, leveraging state space models (SSMs) [28] to causally capture long-range dependencies with linear computational complexity. Recent studies suggest that Mamba-based models can achieve performance comparable to or even better than Transformer-based models. However, their inherent unidirectional sequence modeling in SSMs still poses challenges in achieving an optimal ERF for 2D images (as shown in Fig. 1 (c)). In summary, existing models in MedIR struggle to effectively address the trade-off between a global receptive field and computational efficiency.
@@
In this paper, we propose Restore-RWKV, an efficient and effective model adapted from RWKV for medical image restoration. The original RWKV [35] is designed for handling 1D sequences and has limitations in capturing spatial dependencies in 2D medical images. To address these limitations, we make essential modifications to two key components of RWKV—the attention layer and the token shift layer—to adapt it for processing 2D medical images. (1) We introduce a recurrent WKV (Re-WKV) attention mechanism that effectively captures global dependencies with linear computational complexity. Unlike the causal WKV attention in the original RWKV, which is unidirectional and has a limited receptive field, Re-WKV utilizes a bidirectional attention mechanism to achieve a global receptive field. Additionally, it employs a recurrent attention mechanism that effectively models 2D dependencies across various scan directions. (2) We develop an omnidirectional token shift (Omni-Shift) layer to accurately capture local dependencies. Previous token shift layers only shift adjacent tokens from limited directions using simple interpolation. In contrast, Omni-Shift enhances interaction among neighbors by shifting tokens from all directions and over a large context range through efficient depth-wise convolution. Furthermore, during training, Omni-Shift employs a structural re-parameterization strategy to learn to accurately shift tokens from various context ranges, while maintaining the original structure for efficiency during testing. With these two innovations, the proposed Restore-RWKV effectively models both global and local dependencies in 2D images, achieving the largest global receptive field (as shown in Fig. 1 (d)). Extensive experimental results indicate that our proposed Restore-RWKV can achieve superior performance with good efficiency, serving as a general restoration backbone for various tasks, including MRI image super-resolution, CT image denoising, PET image synthesis, and all-in-one medical image restoration.
@@ 2.1 Medical Image Restoration
High-quality (HQ) medical images, which are essential for accurate disease diagnosis, typically offer superior image quality but come with the drawbacks of longer scanning times or increased radiation exposure. In contrast, low-quality (LQ) images, which reduce scanning time and radiation exposure, often suffer from degraded image quality. As a result, there is significant interest in medical image restoration (MedIR) techniques aimed at recovering HQ medical images from their degraded LQ counterparts. Typical MedIR tasks include PET image synthesis [1, 2, 3, 4, 5, 6, 7, 8], CT image denoising [9, 10, 11, 12, 13, 14, 15], MRI image super-resolution [16, 17, 18, 19, 20], and all-in-one medical image restoration for multi-task MedIR[21, 8, 22, 23]. Recently, deep learning-based models have become the primary approach for addressing MedIR tasks. These methods can be broadly categorized into CNN-based models, Transformer-based models, and Mamba-based models.
CNN-based Models. Over the past decades, CNN-based models [9, 1, 6] have made extraordinary contributions to the field of MedIR due to their strong representation capabilities. A key advantage of CNN-based models for MedIR is their ability to capture local context and build local dependencies through convolution. This allows the neighborhood of a degraded pixel to be used as a reference to aid in its recovery. However, due to the limited kernel size of convolutions, CNN-based methods often suffer from restricted receptive fields and cannot model long-range dependencies. While some approaches have tried to expand the receptive field by designing deeper networks [33, 42] or utilizing the multi-scale U-Net [9, 6] architecture, these improvements are still limited and struggle to capture long-range dependencies effectively.
@@
Mamba-based Models. Mamba [28, 29], a sequence model backbone grounded in state-space models (SSMs), has emerged as a prominent alternative to Transformers. Unlike Transformers, which exhibit quadratic computational complexity, Mamba-based models can model long-range dependencies with linear computational complexity. This capability allows Mamba-based models to potentially capture global dependencies in high-resolution medical images through SSMs without the need for efficient designs like window partitioning. Recent studies have shown that Mamba-based models [32, 15, 30] outperform Transformer-based models in MedIR tasks. However, due to the unidirectional causal attention mechanism of SSM, its receptive field is limited to the range from the first token to the current token in the sequence. As a result, it fails to capture information from subsequent tokens and cannot achieve a global receptive field. While employing a multi-directional scanning strategy [46, 32] can expand the receptive field to encompass a global context, it proves neither efficient nor effective. Consequently, despite Mamba’s ability to model long-range dependencies, it still faces limitations in capturing global dependencies in 2D medical images.
@@ 3 Method
We aim to build an efficient and effective model for medical image restoration based on the Receptance Weighted Key Value (RWKV) [35], termed Restore-RWKV. In Section 3.1, we introduce the network architecture of Restore-RWKV, and in Section 3.2, we present a novel R-RWKV block as its fundamental component for feature extraction. Since the original RWKV [35] is designed for processing 1D sequences, we incorporate two innovations in the R-RWKV block to adapt it for capturing global and local dependencies in 2D images. First, in Section 3.3, we introduce a recurrent WKV (Re-WKV) attention mechanism that effectively builds global dependencies in 2D images with linear computational complexity. Second, in Section 3.4, we introduce an omnidirectional token shift mechanism that accurately captures local context to enhance the model performance. We will elaborate on the details of Restore-RWKV in the following section.
@@ 3.1 Restore-RWKV Architecture
As shown in Fig 2 (a), the proposed Restore-RWKV is a 4-level U-shaped encoder-decoder architecture with 3 times downsampling and upsampling, which enjoys the advantage of capturing image features at different hierarchies and computational efficiency [47]. Given an input low-quality (LQ) image $I_{LQ}\in\mathbb{R}^{H\times W\times 1}$, Restore-RWKV first employs a $3\times 3$ convolutional layer as an input projection to project the input image to a shallow feature $F_{S}\in\mathbb{R}^{H\times W\times C}$, where $H$, $W$, and $C$ denote the height, width, and channel, respectively. Then the shallow feature $F_{S}$ undergoes a 4-level hierarchical encoder-decoder and is transformed into a deep feature $F_{D}\in\mathbb{R}^{H\times W\times 2C}$. Each level of the encoder-decoder comprises $N_{i}$, where $i\in\{1,2,3,4\}$, R-RWKV blocks for feature extraction, with an additional $N_{\text{refinement}}$ R-RWKV blocks used for feature refinement in the final decoder level. For feature downsampling, Restore-RWKV utilizes a $1\times 1$ convolutional layer and a pixel-unshuffle operation [48], reducing the spatial size by half and doubling the channel number. For feature upsampling, Restore-RWKV employs a pixel-shuffle operation [48] and a $1\times 1$ convolutional layer, doubling the spatial size and halving the channel number. To aid in the restoration process, encoder features are concatenated with decoder features via skip connections. The deep feature $F_{D}$ is then projected out to a residual image $I_{R}\in\mathbb{R}^{H\times W\times 1}$ by a $3\times 3$ convolutional layer. Finally, the restored image $\hat{I}_{HQ}$ can be obtained by adding the input $I_{LQ}$ and the residual image $I_{R}$: $\hat{I}_{HQ}=I_{LQ}+I_{R}$.
@@ 3.3 Recurrent WKV Attention
Here, $\operatorname{Bi-WKV}^{(j)}(\cdot)$ denotes the $j$-th Bi-WKV attention. $\operatorname{\Delta}_{dir}(\cdot)$ represents the changing direction operation. In two adjacent Bi-WKV attention operations, we alternately use two different scan directions: horizontal scan (H-Scan) and vertical scan (V-Scan). $wkv^{(j)}$ and $wkv^{(j-1)}$ are attention results of the $j$-th and $(j-1)$-th Bi-WKV, respectively. Note that $wkv^{(0)}$ is the input $V_{s}$. Eq. 9 indicates the current $j$-th Bi-WKV attention takes the previous $(j-1)$-th Bi-WKV attention result as its input value, complementing the $(j-1)$-th attention result from a different scan direction. The final Re-WKV attention result $wkv$ is obtained after applying Bi-WKV attention recurrently $M$ times:
@@
In comparison with Bi-WKV, Re-WKV enhances global token interactions in 2D images by recurrently performing attention along different scan directions. Furthermore, since $M\ll T$ in our implementation, Re-WKV maintains the linear computational complexity as Bi-WKV.
@@ 3.4 Omnidirectional Token Shift
where $\alpha_{i}$ denotes a learnable parameter for scaling the specific branch. $\operatorname{DConv}_{k\times k}(\cdot)$ denotes the efficient depth-wise convolution with kernel size of $k$. The final result of Omni-Shift is a fusion of four branches, with each branch specializing in a specific context range, resulting in an accurate token shift.
@@ 4.1 Dataset
We conduct extensive experiments on various MedIR tasks to demonstrate the performance of our proposed Restore-RWKV model. These include three individual tasks—PET image synthesis, CT image denoising, and MRI image super-resolution—as well as an all-in-one medical image restoration task that combines the aforementioned three tasks for multi-task learning. For each individual MedIR task, we prepare two types of datasets: a base dataset for model training and in-distribution testing, and another dataset for out-of-distribution testing. The datasets corresponding to each task are described below.
PET Image Synthesis. (1) PolarStar m660 Dataset. The PolarStar m660 dataset is a private dataset utilized for both model training and in-distribution testing. It contains HQ PET images collected from 159 patients using the PolarStar m660 PET/CT system in list mode, with an average injection dose of 293MBq 18F-FDG. These 159 HQ PET images were divided into 120 for training, 10 for validation, and 29 for testing. LQ PET images are generated through random list mode decimation with a dose reduction factor (DRF) of 12. Both HQ and LQ PET images are reconstructed using the standard OSEM [50] method. Each PET image has 3D shapes of 400$\times$192$\times$192 and is divided into 192 2D slices sized 400$\times$192. Slices containing only air are excluded. Consequently, the PolarStar m660 dataset contains 8350 2D PET images for training, 684 images for validation, and 2044 images for testing. (2) PolarStar Flight Dataset. The PolarStar Flight dataset is another private dataset for out-of-distribution testing. It contains HQ PET images collected from 10 patients using the PolarStar Flight PET/CT system in list mode, with an average injection dose of 293MBq 18F-FDG. The DRF for the LQ images is set to 4. The data preparation process is identical to that of the PolarStar m660 dataset. In total, the PolarStar Flight dataset contains 674 image pairs, each with a size of 400$\times$192.
CT Image Denoising. (1) AAPM Dataset. The publicly available AAPM dataset [51] is employed for both model training and in-distribution testing. It comprises paired standard-dose HQ CT images and quarter-dose LQ CT images of the abdomen. These images are collected from 10 patients and are divided into 8 patients for training, 1 for validation, and 1 for testing. Each 3D CT image has a dimension of 512$\times$512×$\times n$. Consequently, we extract 2D slices sized 512$\times$512 and obtain 2039 images for training, 128 for validation, and 211 for testing. (2) LDCT Dataset. The publicly available LDCT dataset [52] is used for out-of-distribution testing. We randomly select 5 paired standard-dose HQ CT images and quarter-dose LQ CT images of the abdomen. After extracting 2D slices, the LDCT dataset contains 747 image pairs, each with dimensions of 512$\times$512.
@@
Table 1: PET image synthesis results. The best result is marked in bold, and the second-best result is underlined.
PolarStar m660 PolarStar Flight
Method Params (M) FLOPs (G) PSNR↑ SSIM↑ RMSE↓ PSNR↑ SSIM↑ RMSE↓
Xiang’s 0.23 3.77 35.93 0.9167 0.0980 33.85 0.9092 0.1320
Xu’s 0.52 1.11 36.44 0.9376 0.0982 34.14 0.9160 0.1275
DCNN 1.33 21.86 36.27 0.9243 0.0954 33.96 0.9135 0.1303
CycleWGAN 1.00 16.42 36.62 0.9290 0.0910 34.12 0.9132 0.1279
DCITN 0.08 1.24 36.09 0.9285 0.0970 33.68 0.9064 0.1346
ARGAN 31.14 8.48 36.73 0.9406 0.0902 34.13 0.9191 0.1279
DRMC 0.62 9.92 36.00 0.9352 0.0998 33.62 0.9081 0.1356
[…]
Restore-RWKV 27.91 37.46 37.33 0.9474 0.0852 34.70 0.9256 0.1212
Figure 5: Visual comparison of different methods in PET image synthesis. The zoomed-in rectangular region is recommended for a better comparison.
Figure 6: Performance comparison on (a) clinical ROIs and (b) low-dose PET images with different DRFs.
All-in-One Medical Image Restoration. The objective of all-in-one medical image restoration [22] is to train a universal model capable of handling multiple MedIR tasks. Therefore, the dataset used in this study combines datasets from the aforementioned three individual tasks: the PolarStar m660 dataset for PET image synthesis, the AAPM dataset for CT image denoising, and the IXI dataset for MRI image super-resolution. The training, validation, and testing splits for each dataset remain consistent with their original partitions used in the individual tasks, but they are combined into a unified dataset for this study.